# Collisional Baryon-Dominated Dwarf Galaxies: A New Probe of Bursty Feedback and Dark Matter Physics

### v1

@@ Abstract
High-velocity collisions between gas-rich ultra-diffuse galaxies present a promising formation channel for baryon-dominated dwarf galaxies (BDDGs). Using hydrodynamical simulations, we show that the progenitors’ baryonic binding energy, $|E_{\rm bind}|$, critically controls the outcome. Repeated potential fluctuations, e.g., from bursty feedback, inject energy and reduce $|E_{\rm bind}|$ by $\approx 15\%$, yielding fewer but substantially more massive BDDGs. By contrast, elastic self-interacting dark matter produces comparable cores without lowering $|E_{\rm bind}|$, resulting in negligible effect. This provides a novel way to distinguish between two leading galactic core formation channels, i.e., the baryon feedback and elastic dark matter self-interaction. Among 15 paired simulation runs, 13 show higher BDDG masses in the weakened-binding case, and about two thirds exhibit $100\%$ mass enhancements. The simulations also predict systematically lower gas fractions due to sustained post-collision star formation, yielding a clean observational signature. Upcoming wide-field imaging (CSST, LSST), HI surveys (FAST), and kinematic follow-up will be crucial to test this scenario.
Within the standard cosmological model, the cosmic baryon fraction is only one-fifth that of the dark matter, and galaxies are expected to reside in more massive dark matter halos. As a result of galaxy formation governed by complex processes, the dark matter content of typical dwarf galaxies is expected to exceed their stellar mass by roughly two orders of magnitude [1, 2, 3, 4].
@@
Notably, Guo et al. [7] identified $19$ DMDGs from a sample of $324$ SDSS dwarfs, with $14$ located well beyond the virial radii of nearby groups or clusters, where environmental stripping and recent interactions are unlikely. In a related context, tidal dwarf galaxies (TDGs) are also baryon dominated, forming from the tidal debris of interacting galaxies [17, 18, 19, 20]. Kaviraj et al. [19] estimated that about 66% of dwarfs in nearby clusters may have a tidal origin, while Zaragoza-Cardiel et al. [20] identified $39$ TDGs in the local universe that appear to have escaped their host halos. These findings suggest that tidal interactions play a crucial role in producing baryon-dominated galaxies, with galaxy collisions representing the most extreme and favorable conditions for their formation. In this work, we adopt the term baryon-dominated dwarf galaxies (BDDGs) to broadly denote systems in which baryons outweigh dark matter, as expected in collision-driven formation.
@@
In high-speed dwarf–dwarf collisions, ram pressure and shocks can exert forces that separate gas from dark matter, allowing the gas to reassemble into bound, baryon-dominated remnants. Several recent studies have simulated such collisions and demonstrated that gas-rich dwarfs can, under specific conditions, produce remnants with low dark matter content and diffuse morphologies [e.g., 21, 22, 16]. These simulations have examined how collision velocity, disk structure, and initial configuration affect the formation of BDDGs. However, it remains unclear what physical quantity critically controls the mass and yield of BDDGs.
In this work, we identify the baryonic binding energy of the progenitors, $|E_{\rm bind}|$, as the relevant control parameter, directly linking energy injection in dwarf halos to BDDG yields. We employ controlled hydrodynamical simulations of gas-rich ultra-diffuse galaxy (UDG) collisions to assess the potential of these systems as probes of energy injection in dwarf halos, positioning BDDG observations a new observational window for testing the underlying physics. Intriguingly, both baryonic feedback [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] and self-interacting dark matter (SIDM) [35, 36, 37, 38, 39, 40, 41] have been proposed as mechanisms for generating cores in halos, but their consequences diverge remarkably in collisional events. Future BDDG studies can thus inform both galaxy formation and dark sector physics.
Weakening gravitational binding by energy injection. To unbind baryons from halos more efficiently, we need processes that reduce the gravitational binding of the entire system. Although energy injection into halos can weaken the binding of their baryonic component, most baryonic processes, such as gas cooling, release their energy radiatively, thereby dissipating rather than adding energy to the gravitationally bound system. In contrast, repeated potential fluctuations can heat dark matter in the inner regions, reducing the gravitational binding after relaxation and flattening the inner halo density profile. Such fluctuations arise naturally from bursty stellar feedback in dwarf galaxies [28, 42]. In novel dark matter models, recurrent collapses of solitonic cores can trigger bosenova-like outbursts that drive similar fluctuations [43, 44]. Scattering among dark matter particles may also be exothermic, imparting kinetic energy to the final-state particles while keeping them bound to the halo [45, 46].
@@
As a quantitative illustration, we model the evolution of a halo density profile under energy injection, following the method of Ref. [28], which has been widely applied in various contexts [28, 47, 48, 49]. For demonstration, we present results for a spherical, isotropic NFW halo with a mass of $1.5\times 10^{10}~\rm M_{\odot}$ and a concentration parameter of $c=14$, corresponding to the 1a and 1b benchmarks in Table 1. The baryon content is assumed to be a gas disk which taks spherically averaged pseudo-isothermal profile, $\rho_{\rm gas}=\rho_{0}\left(1+(r/r_{g})^{2}\right)^{-1}$, where $r_{g}=2~\rm kpc$ and $\rho_{0}=2.46\times 10^{6}~\rm M_{\odot}~\rm kpc^{-3}$ [50, 16]. The disk height is set to $0.2~\rm kpc$.
@@
Figure 1 shows the evolution of the density profile (left), gravitational potential (middle), and $|E_{\rm bind}|$ (right) in 2 Gyr. We compute $E_{\rm bind}$ by integrating the gas density over the total gravitational potential ($\Psi_{\rm tot}$), $E_{\rm bind}=4\pi\int drr^{2}\Psi_{\rm tot}\rho_{\rm gas}$. The results illustrate that impulsive feedback heats the inner orbits, leading to an expansion of centrally cored density profile, a shallowing of the central potential, and a reduction in the central baryonic binding energy. By $2$ Gyr, $|E_{\rm bind}|$ has decreased by over 10%. We will use collision simulations to show that even this modest reduction can substantially enhance the BDDG formation.
We also contrast this with the effect of SIDM, a distinct core formation mechanism. To this end, we simulate the same initial condition as in the 1a and 1b benchmarks (Table 1), considering only gravity and using the Gadget-2-based SIDM module developed in Yang et al. [51, 52], adopting a cross section of $\sigma/m=20~\rm{cm^{2}g^{-1}}$.
In Fig. 1, the evolved density profile at $2~\rm Gyr$ is shown in gray. While the SIDM halo develops a core of comparable size to that formed by feedback, its gravitational potential remains deep, and $|E_{\rm bind}|$ slightly increases rather than decreases. This suggests that baryonic binding, if quantified observationally, could serve as a diagnostic to distinguish between the two core formation mechanisms.
Figure 2: Formation of BDDGs from the collision of gas-rich dwarf galaxies. The initial setup (left) has two progenitors approaching at a close separation. They collide within 0.2 Gyr (middle), displacing gas from the halo centers and triggering efficient star formation. By 5 Gyr (right), the expelled baryons have collapsed into several BDDGs that are far from their progenitors. The surface density distributions of dark matter (top), gas (middle), and stars (bottom) are presented at the corresponding snapshots for the 2a ($\gamma=1$) benchmark.
Simulation setup. To illustrate the impact of energy injection into halos in a model-independent manner, we adopt the following NFW-like extension for the halo density profile [53, 10]
We construct two contrasting sets of initial conditions, one with $\gamma=1$ (cuspy) and the other with $\gamma=0.1$ (cored, with weaker binding), while keeping the total halo mass and concentration fixed. The simulated halos span masses of $1-2\times 10^{10}~\rm M_{\odot}$ and concentrations in the range 4144-14. Assuming these halos host gas-rich UDGs, we model the baryonic content with a gas-only disk, as described previously, and vary the disk height between $0.15$ and $0.35~\rm kpc$. This maximizes the effectiveness of BDDG formation, as stars behave as a collisionless component during the collision and cannot be efficiently expelled from the progenitors. The gas-to-halo mass ratio is set to $0.1$ or $0.22$, with gas scale radii $r_{g}$ in the range 1.53.51.5-3.5 kpc. These choices are motivated by observations of gas-rich UDGs, where internal turbulence and stellar feedback maintain gas densities below the threshold for molecular hydrogen formation, $n_{g}\lesssim 1~\mathrm{cm}^{-3}\approx 3\times 10^{7}~M_{\odot}\mathrm{kpc}^{-3}$, thereby suppressing star formation [54, 42]. The population and properties of gas-rich UDGs have attracted considerable attention and been extensively explored in recent studies [55, 56, 42, 57, 58, 59].
Equilibrium initial conditions are generated using DICE [60], with the gas disk stabilized by enforcing a Toomre $Q_{\rm gas}>1.5$ [61, 62]. The gas is initialized at $T=10^{4}$ K and metallicity $Z=0.1Z_{\odot}$. We evolve the collision simulations using Gadget-4 [63], incorporating radiative cooling of primordial gas and star formation based on the multiphase ISM model of Springel & Hernquist (2003) [64]. Stellar feedback is modeled as an effective pressure that prevents the gas from collapsing too quickly. To obtain the SIDM prediction, we set up a hydrodynamical CDM simulation using progenitors taken from the isolated SIDM simulation (Gadget-2) snapshot at 2 Gyr. We verified that switching off SIDM after this initialization yields a stationary profile in isolation. Further details are provided in the Supplementary Material.
Our collision simulations build upon findings in previous studies, where dependencies of simulation outcomes on the collision velocities, progenitor halo mass and concentration, gas distribution, and numerical resolution were systematically explored. We adopt simulation setups similar to those in Refs. [21, 22, 16], with relative collision velocities ($v_{r}$) ranging from $300-600~\rm km/s$. Our primary choice of $v_{r}=400~\rm km/s$ lies within the preferred range reported in the literature. We also perform simulations with $v_{r}=280$, $450$ and $500~\rm km/s$. The two colliding progenitors are initialized with identical conditions, placed at a separation of $60$ kpc along their relative velocity vector and with an impact parameter of $2$ kpc. We focus on contrasting paired samples with $\gamma=1$ and $\gamma=0.1$, testing them under a range of conditions. Table 1 summarizes the representative benchmarks adopted in our simulations. Full details of the initial conditions and parameter setup are provided in the Supplementary Material.
Table 1: Representative simulation benchmarks. The full simulation parameters and numerical results are provided in the Supplemental Material.
@@
Collisional formation of BDDGs. Figure 2 shows simulation snapshots of the surface densities of dark matter (top), gas (middle), and stars (bottom) for the benchmark 2a ($\gamma=1$) in Table 1. Four stages of the collision are displayed. The first snapshot depicts the initial setup, with two progenitors approaching each other at close separation so that the simulation outcome primarily concerns the formation of BDDGs rather than the internal dynamics of the progenitors. The systems collide within 0.2 Gyr, displacing their gas from the halo centers and triggering efficient star formation. By 5 Gyr, the expelled baryons have already collapsed into several BDDGs and moved far from their progenitors.
@@
With $v_{r}=400~\rm km/s$, most of the progenitor gas is expelled. In the cuspy case ($\gamma=1$), the steeper dark-matter potential exerts stronger tidal “winds,” causing the expelled gas to fragment into several small, low-mass condensations. In contrast, in the weakened-binding case ($\gamma=0.1$), the debris remains in contact more easily, allowing it to coalesce into fewer but more massive BDDGs.
@@
The $\gamma=0.1$ model develops broad, contiguous compressive regions around the origin across all separations, consistent with debris that readily coalesces into fewer, more massive clumps. A shallower inner slope renders the central potential more nearly harmonic, weakens shear, and suppresses stretching along the collision axis. By contrast, the cuspy $\gamma=1$ case exhibits only small, fragmented compressive islands around shear-dominated zones, consistent with the formation of multiple low-mass condensations. The SIDM core is encircled by a ring of stretching tides that inhibits gas confinement. The central compressive region still facilitates gas condensation, but the formation of BDDG is significantly suppressed compared with the $\gamma=0.1$ scenario.
@@
In a weakened gravitational binding scenario, gas is expected to escape more readily during collisions. To illustrate this, we reduce the collision velocity of the 3a and 3b benchmarks to $280~\mathrm{km/s}$ (see Table 1 for setup and results). The escaping gas mass reaches $3.8\times 10^{9}~M_{\odot}$ in the $\gamma=0.1$ case, compared to $3.3\times 10^{9}~M_{\odot}$ for $\gamma=1$. Both the stellar and gas masses bound to the resulting BDDGs are substantially higher in the $\gamma=0.1$ scenario, with enhancements of $240\%$ and $34\%$, respectively. The most massive BDDG formed in our simulations has a stellar mass of $1.68\times 10^{9}~M_{\odot}$, obtained in the $\gamma=0.1$, $v_{r}=280~\mathrm{km/s}$ run. These results demonstrate that reduced baryonic binding favors the formation of more massive BDDGs.
@@
Figure 4: Characteristics of the most massive BDDGs in 15 paired simulations. The bar chart summarizes results from contrasting simulations with inner slopes $\gamma=0.1$ and $\gamma=1$. Bars are arranged in ascending order of the BDDG mass ratio, $M_{b,\gamma=0.1}/M_{b,\gamma=1}$. The bar width scales with the relative change in binding energy, $(E_{{\rm bind},\gamma=1}-E_{{\rm bind},\gamma=0.1})/E_{{\rm bind},\gamma=1}$, while the color indicates the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. In contrast, the increase in the BDDG masses is predominantly positive. In 9 out of the 15 simulations, this increase exceeds 100%, as reflected by the bars that rise above the line representing $\Delta M_{b}/M_{b,\gamma=1}=1$.
@@
Aside from the examples shown, we simulated 15 pairs of dwarf collisions with varied setups. Figure 4 presents the relative BDDG mass enhancements in the $\gamma=0.1$ and $\gamma=1$ cases, with bar widths proportional to the reduction in baryonic binding energy and colors indicating the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. We find that 13 of the 15 simulations yield an increase in BDDG mass, with 9 of these cases showing an enhancement greater than 100%. The two exceptions occur for halos with extremely low concentrations ($c=4$) and $v_{r}=450~\mathrm{km/s}$, conditions that reduce the contrast between the two scenarios. Interestingly, the gas fractions in the $\gamma=0.1$ runs are systematically lower than in their $\gamma=1$ counterparts, demonstrating more sustained star formation following the collision. In this figure, we illustrate results considering only the most massive BDDGs. Our findings remain largely unchanged when considering the sum of multiple BDDG masses. The Supplementary Material presents analogous results that include the total BDDG mass as well as the case with the gas cooling rate reduced by half. Additional details on post-collision star formation are also provided.

### v2

# Dark-Matter-Deficient Galaxies from Collisions: A New Probe of Bursty Feedback and Dark Matter Physics

@@ Abstract
High-velocity collisions between gas-rich ultra-diffuse galaxies present a promising formation channel for dark-matter-deficient galaxies (DMDGs). Using hydrodynamical simulations, we show that the progenitors’ baryonic binding energy, $|E_{\rm bind}|$, critically controls the outcome. Repeated potential fluctuations, e.g., from bursty feedback, inject energy and reduce $|E_{\rm bind}|$ by $\approx 15\%$, yielding fewer but substantially more massive DMDGs. By contrast, elastic self-interacting dark matter (SIDM) produces comparable cores without lowering $|E_{\rm bind}|$, resulting in negligible effect. This differs from the other DMDG formation scenario, where SIDM-induced cores enhance dark matter tidal stripping while keeping baryons compact and resilient to tidal effects. The contrasting roles of SIDM help distinguish feedback-formed halo cores from those created by SIDM. Among 15 paired simulation runs, 13 show higher DMDG masses in the weakened-binding case, and about two thirds exhibit $100\%$ mass enhancements. The simulations also predict systematically lower gas fractions due to sustained post-collision star formation, yielding a clean observational signature. Upcoming wide-field imaging (CSST, LSST), HI surveys (FAST), and kinematic follow-up will be crucial to test this scenario.
show]yangdn@pmo.ac.cn
show]yzfan@pmo.ac.cn
1 Introduction
Within the standard cosmological model, the cosmic baryon fraction is only one-fifth that of the dark matter, and galaxies are expected to reside in more massive dark matter halos. As a result of galaxy formation governed by complex processes, the dark matter content of typical dwarf galaxies is expected to exceed their stellar mass by roughly two orders of magnitude (A. Dekel & J. Silk, 1986; P. F. Hopkins et al., 2014; A. Vale & J. P. Ostriker, 2004; P. S. Behroozi et al., 2013).
@@ 1 Introduction
Existing cosmological simulation studies have identified DMDGs as tidally stripped satellites (J. Moreno et al., 2022; Y. Jing et al., 2019). In this scenario, controlled simulations further map the orbital conditions and galaxy-halo configurations required to form such systems, offering a way to test models of elastic self-interacting dark matter (SIDM) that produce cored halos without making baryons too diffuse (D. Yang et al., 2020). However, the identification of DMDGs in isolation remains challenging, hindered by the competing requirements for both high resolution and a large field of view. Recently, J. Lee et al. (2024) used the TNG100-1 simulation (D. Nelson et al., 2019) to search for gas-rich dwarf pairs whose orbital parameters satisfy requirements for producing DMDGs near massive hosts, finding $\sim 10$ such collisions in a $(100\ {\rm Mpc})^{3}$ volume over $z\simeq 3\rightarrow 0$. The mass and spatial resolution of TNG100 are insufficient to follow the hydrodynamics required to produce such systems, but the implied event rate broadly agrees with the handful of field DMDGs reported by Q. Guo et al. (2019), suggesting that dwarf-dwarf collisions provide a cosmologically plausible channel for generating observable DMDGs. Though how collision velocity, disk structure, and initial configuration affect the DMDG formation has been examined in some recent simulations (E.-j. Shin et al., 2020; J. Lee et al., 2024), it remains unclear what physical quantity critically controls the mass and yield of DMDGs.
In this work, we identify the baryonic binding energy of the progenitors, $|E_{\rm bind}|$, as the relevant control parameter, directly linking energy injection in dwarf halos to DMDG yields. We employ controlled hydrodynamical simulations of gas-rich ultra-diffuse galaxy (UDG) collisions to assess the potential of these systems as probes of energy injection in dwarf halos, positioning DMDG observations a new observational window for testing the underlying physics. Intriguingly, both baryonic feedback (J. F. Navarro et al., 1996; S. Mashchenko et al., 2006; F. Governato et al., 2010) and self-interacting dark matter (SIDM) (D. N. Spergel & P. J. Steinhardt, 2000; S. Tulin & H.-B. Yu, 2018)have been proposed as mechanisms for generating cores in halos, but their consequences diverge remarkably in collisional events. Future DMDG studies can thus inform both galaxy formation and dark sector physics.
2 Weakening gravitational binding by energy injection.
To unbind baryons from halos more efficiently, we need processes that reduce the gravitational binding of the entire system. Although energy injection into halos can weaken the binding of their baryonic component, most baryonic processes, such as gas cooling, release their energy radiatively, thereby dissipating rather than adding energy to the gravitationally bound system. In contrast, repeated potential fluctuations can heat dark matter in the inner regions, reducing the gravitational binding after relaxation and flattening the inner halo density profile. Such fluctuations arise naturally from bursty stellar feedback in dwarf galaxies (A. Pontzen & F. Governato, 2012; T. K. Chan et al., 2018; J. Freundlich et al., 2020; Z. Li et al., 2023). In novel dark matter models, recurrent collapses of solitonic cores can trigger bosenova-like outbursts that drive similar fluctuations (P. J. Fox et al., 2023; H. Koo, 2025). Scattering among dark matter particles may also be exothermic, imparting kinetic energy to the final-state particles while keeping them bound to the halo (M. Vogelsberger et al., 2019; M. V. Medvedev, 2014).
@@
As a quantitative illustration, we model the evolution of a halo density profile under energy injection, following the method of A. Pontzen & F. Governato (2012), which has been widely applied in various contexts (A. Pontzen & F. Governato, 2012; A. Di Cintio et al., 2014; S. Peirani et al., 2017; J. F. Acevedo et al., 2024). For demonstration, we present results for a spherical, isotropic NFW halo with a mass of $1.5\times 10^{10}~\rm M_{\odot}$ and a concentration parameter of $c=14$, corresponding to the 1a and 1b benchmarks in Table 1. For each dark matter particle, we compute the total energy $E_{0}$ and specific angular momentum $j$. A sequence of impulsive potential fluctuations is modeled by rescaling the potential by a small fraction, $\Delta V/V_{0}=0.14$, at each iteration. The energy of each particle is updated using the phase-averaged kick at fixed $j$
$\Delta E=\frac{2nE_{0}}{(2+n)^{2}}\,\bigg(\frac{\Delta V}{V_{0}}\bigg)^{2},$ (1)
$p(r;E,j)$ $\propto$ $\frac{1}{\sqrt{E-V_{\rm eff}(r,j)}},$ (2)
$V_{\rm eff}(r,j)$ $=$ $V(r)+\frac{j^{2}}{2r^{2}},$ (3)
over all dark matter particles. The baryon content is assumed to be a gas disk which takes spherically averaged pseudo-isothermal profile, $\rho_{\rm gas}=\rho_{0}\left(1+(r/r_{g})^{2}\right)^{-1}$, where $r_{g}=2~\rm kpc$ and $\rho_{0}=2.46\times 10^{6}~\rm M_{\odot}~\rm kpc^{-3}$ (S. Kurapati et al., 2020; J. Lee et al., 2024). The disk height is set to $0.2~\rm kpc$.
@@
Figure 1 shows the evolution of the density profile (left), gravitational potential (middle), and $|E_{\rm bind}|$ (right) in 2 Gyr. We compute $E_{\rm bind}$ by integrating the gas density over the total gravitational potential ($\Psi_{\rm tot}$), $E_{\rm bind}=4\pi\int drr^{2}\Psi_{\rm tot}\rho_{\rm gas}$. The results illustrate that impulsive feedback heats the inner orbits, leading to an expansion of centrally cored density profile, a shallowing of the central potential, and a reduction in the central baryonic binding energy. By $2$ Gyr, $|E_{\rm bind}|$ has decreased by over 10%. We will use collision simulations to show that even this modest reduction can substantially enhance the DMDG formation.
Table 1: Representative simulation benchmarks.
@@
Note. — The full simulation parameters and numerical results are provided in the Materials and Methods section.
We also contrast this with the effect of SIDM, a distinct core formation mechanism. To this end, we simulate the same initial condition as in the 1a and 1b benchmarks (Table 1), considering only gravity and using the Gadget-2-based SIDM module developed in D. Yang & H.-B. Yu (2022) and D. Yang et al. (2023), adopting a cross section of $\sigma/m=20~\rm{cm^{2}g^{-1}}$. In Figure 1, the evolved density profile at $2~\rm Gyr$ is shown in gray. While the SIDM halo develops a core of comparable size to that formed by feedback, its gravitational potential remains deep, and $|E_{\rm bind}|$ slightly increases rather than decreases. This suggests that baryonic binding, if quantified observationally, could serve as a diagnostic to distinguish between the two core formation mechanisms.
Table 1 summarizes the configurations and results of representative simulations. The labels $\gamma=0.1$ and $\gamma=1$ correspond to simulations with and without weakened baryonic gravitational binding. We model the density profiles using an extended NFW profile with an inner slope controlled by the parameter $\gamma$ (see Materials and Methods for details).
3 Simulation setup
To illustrate the impact of energy injection into halos in a model-independent manner, we adopt the following NFW-like extension for the halo density profile (A. W. Graham et al., 2006; G. Ogiya, 2018)
We construct two contrasting sets of initial conditions, one with $\gamma=1$ (cuspy) and the other with $\gamma=0.1$ (cored, with weaker binding), while keeping the total halo mass and concentration fixed. The simulated halos span masses of $1-2\times 10^{10}~\rm M_{\odot}$ and concentrations in the range 4144-14. Assuming these halos host gas-rich UDGs, we model the baryonic content with a gas-only disk, as described previously, and vary the disk height between $0.15$ and $0.35~\rm kpc$. This maximizes the effectiveness of DMDG formation, as stars behave as a collisionless component during the collision and cannot be efficiently expelled from the progenitors. The gas-to-halo mass ratio is set to $0.1$ or $0.22$, with gas scale radii $r_{g}$ in the range 1.53.51.5-3.5 kpc. The pericentric distance is assumed to be $2\,\rm kpc$. These choices are motivated by observations of gas-rich UDGs, where internal turbulence and stellar feedback maintain gas densities below the threshold for molecular hydrogen formation, $n_{g}\lesssim 1~\mathrm{cm}^{-3}\approx 3\times 10^{7}~M_{\odot}\mathrm{kpc}^{-3}$, thereby suppressing star formation (A. Di Cintio et al., 2017; T. K. Chan et al., 2018). The population and properties of gas-rich UDGs have attracted considerable attention and been extensively explored in recent studies (E. Papastergis et al., 2017; F. Jiang et al., 2019; T. K. Chan et al., 2018; M. G. Jones et al., 2018; P. E. Mancera Piña et al., 2020; D. Kong et al., 2022).
Equilibrium initial conditions are generated using DICE (V. Perret et al., 2014), with the gas disk stabilized by enforcing a Toomre $Q_{\rm gas}>1.5$ (A. Toomre, 1964; J. Binney & S. Tremaine, 2008; D. Williamson et al., 2016). The gas is initialized at $T=10^{4}$ K and metallicity $Z=0.1Z_{\odot}$. We evolve the collision simulations using Gadget-4 (V. Springel et al., 2021), incorporating radiative cooling of primordial gas and star formation based on the multiphase ISM model of V. Springel & L. Hernquist (2003). Stellar feedback is modeled as an effective pressure that prevents the gas from collapsing too quickly. To obtain the SIDM prediction, we set up a hydrodynamical CDM simulation using progenitors taken from the isolated SIDM simulation (Gadget-2) snapshot at 2 Gyr. We verified that switching off SIDM after this initialization yields a stationary profile in isolation. Further details are provided in the Supplementary Material.
Our collision simulations build upon findings in previous studies, where dependencies of simulation outcomes on the collision velocities, progenitor halo mass and concentration, gas distribution, and numerical resolution were systematically explored. We adopt simulation setups similar to those in E.-j. Shin et al. (2020) and J. Lee et al. (2024), with relative collision velocities ($v_{r}$) ranging from $300-600~\rm km/s$. Our primary choice of $v_{r}=400~\rm km/s$ lies within the preferred range reported in the literature. We also perform simulations with $v_{r}=280$, $450$ and $500~\rm km/s$. The two colliding progenitors are initialized with identical conditions, placed at a separation of $60$ kpc along their relative velocity vector and with an impact parameter of $2$ kpc. We focus on contrasting paired samples with $\gamma=1$ and $\gamma=0.1$, testing them under a range of conditions. Table 2 summarizes the initial conditions and parameter setup in our simulations.
@@ 3 Simulation setup
Figure 2: Formation of DMDGs from the collision of gas-rich dwarf galaxies. The initial setup (left) has two progenitors approaching at a close separation. They collide within 0.2 Gyr (middle), displacing gas from the halo centers and triggering efficient star formation. By 5 Gyr (right), the expelled baryons have collapsed into several DMDGs that are far from their progenitors. The surface density distributions of dark matter (top), gas (middle), and stars (bottom) are presented at the corresponding snapshots for the 2a ($\gamma=1$) benchmark.
4 Collisional formation of DMDGs.
Figure 2 shows simulation snapshots of the surface densities of dark matter (top), gas (middle), and stars (bottom) for the benchmark 2a ($\gamma=1$) in Table 1. Four stages of the collision are displayed. The first snapshot depicts the initial setup, with two progenitors approaching each other at close separation so that the simulation outcome primarily concerns the formation of DMDGs rather than the internal dynamics of the progenitors. The systems collide within 0.2 Gyr, displacing their gas from the halo centers and triggering efficient star formation. By 5 Gyr, the expelled baryons have already collapsed into several DMDGs and moved far from their progenitors.
@@
With $v_{r}=400~\rm km/s$, most of the progenitor gas is expelled. In the cuspy case ($\gamma=1$), the steeper dark-matter potential exerts stronger tidal “winds,” causing the expelled gas to fragment into several small, low-mass condensations. In contrast, in the weakened-binding case ($\gamma=0.1$), the debris remains in contact more easily, allowing it to coalesce into fewer but more massive DMDGs.
@@
The $\gamma=0.1$ model develops broad, contiguous compressive regions around the origin across all separations, consistent with debris that readily coalesces into fewer, more massive clumps. A shallower inner slope renders the central potential more nearly harmonic, weakens shear, and suppresses stretching along the collision axis. By contrast, the cuspy $\gamma=1$ case exhibits only small, fragmented compressive islands around shear-dominated zones, consistent with the formation of multiple low-mass condensations. The SIDM core is encircled by a ring of stretching tides that inhibits gas confinement. The central compressive region still facilitates gas condensation, but the formation of DMDG is significantly suppressed compared with the $\gamma=0.1$ scenario.
@@
In a weakened gravitational binding scenario, gas is expected to escape more readily during collisions. To illustrate this, we reduce the collision velocity of the 3a and 3b benchmarks to $280~\mathrm{km/s}$ (see Table 1 for setup and results). The escaping gas mass reaches $3.8\times 10^{9}~M_{\odot}$ in the $\gamma=0.1$ case, compared to $3.3\times 10^{9}~M_{\odot}$ for $\gamma=1$. Both the stellar and gas masses bound to the resulting DMDGs are substantially higher in the $\gamma=0.1$ scenario, with enhancements of $240\%$ and $34\%$, respectively. The most massive DMDG formed in our simulations has a stellar mass of $1.68\times 10^{9}~M_{\odot}$, obtained in the $\gamma=0.1$, $v_{r}=280~\mathrm{km/s}$ run. These results demonstrate that reduced baryonic binding favors the formation of more massive DMDGs.
@@
Figure 4: Characteristics of the most massive DMDGs in 15 paired simulations. The bar chart summarizes results from contrasting simulations with inner slopes $\gamma=0.1$ and $\gamma=1$. Bars are arranged in ascending order of the DMDG mass ratio, $M_{b,\gamma=0.1}/M_{b,\gamma=1}$. The bar width scales with the relative change in binding energy, $(E_{{\rm bind},\gamma=1}-E_{{\rm bind},\gamma=0.1})/E_{{\rm bind},\gamma=1}$, while the color indicates the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. In contrast, the increase in the DMDG masses is predominantly positive. In 9 out of the 15 simulations, this increase exceeds 100%, as reflected by the bars that rise above the line representing $\Delta M_{b}/M_{b,\gamma=1}=1$.
@@
Aside from the examples shown, we simulated 15 pairs of dwarf collisions with varied setups. Figure 4 presents the relative DMDG mass enhancements in the $\gamma=0.1$ and $\gamma=1$ cases, with bar widths proportional to the reduction in baryonic binding energy and colors indicating the gas fraction, $f=M_{\rm gas}/(M_{\rm gas}+M_{\rm stars})$. We find that 13 of the 15 simulations yield an increase in DMDG mass, with 9 of these cases showing an enhancement greater than 100%. The two exceptions occur for halos with extremely low concentrations ($c=4$) and $v_{r}=450~\mathrm{km/s}$, conditions that reduce the contrast between the two scenarios. Interestingly, the gas fractions in the $\gamma=0.1$ runs are systematically lower than in their $\gamma=1$ counterparts, demonstrating more sustained star formation following the collision. In this figure, we illustrate results considering only the most massive DMDGs. Our findings remain largely unchanged when considering the sum of multiple DMDG masses. The Materials and Methods section presents analogous results that include the total DMDG mass as well as the case with the gas cooling rate reduced by half. Additional details on post-collision star formation are also provided.
5 Discussion and Conclusion